OpenStudy (anonymous):

A trapezoid midsegment measures 6. One of the bases measures 10. What is the measure of the other base?

OpenStudy (anonymous):

Im failing math and really need help! Thanks!

OpenStudy (anonymous):

10 + b ------ = 6 2 10 + b = 12 b = ? ?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

what does b equal?

OpenStudy (anonymous):

if 10+b=12 what does b equal?

OpenStudy (anonymous):

idk thats whats i need lol like one of the bases measure 10 and the midsegment measures 10

OpenStudy (anonymous):

sryy i'm not gud at explaining things

OpenStudy (anonymous):

2 8 14 16

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

np